# Magic-Boost: Boost 3D Generation with Mutli-View Conditioned Diffusion

### v1

@@ 3.3 Refinement with SDS Optimization
As shown in Fig. 2, We build our test pipeline with Instant3D [13], a two-stage feed-forward generation method which firstly generates four multi-view images with finetuned 2D diffusion networks and then lift the multi-view images to 3D model utilizing large reconstruction model. However, due to limited resolution and local inconsistency, the generated results are still in low quality without detailed texture and complicated geometry. Leveraging the generated pseudo multi-view as input, we propose to adopt SDS optimization with small noise level to further enhance the coarse generation results. Benefit from the high-fidelity geometry and texture information provided by the pesudo multi-view inputs, our model is capable of generating highly consistent novel views and providing precise SDS guidance to enhance the coarse generated results within a short time period ($\sim 15$min).
@@
We further introduce an Anchor Iterative Update loss to alleviate the over-saturation problem of SDS optimization. Specifically, we draw inspiration from the recent image editing methods [9, 22] that edit the NeRF by alternatively rendering dataset image, updating the dataset and then supervising the NeRF reconstruction with the updated dataset images. In our generation task, we could also regard the input pesudo multi-view inputs as our initial anchor datasets and adopt a similar update strategy by first rendering anchor view image, perturbing the image with random noise and then apply a multi-step denoising process with the proposed multi-view condition diffusion model. However, as the denoising process is guided by the multi-view inputs themselves, simply adopt such as process leading to minor refinement on the input anchor images. To address this problem, while updating certain anchor view $v_{1}$, we drop it in the condition inputs for the multi-view diffusion models, and denoise leveraging other views $v_{0},v_{2},v_{3}$. We found this works well and achieve accurate refinement on the local details of the anchor image, while preserving the 3D consistency. The refined anchor image is then used to supervise the generation with MSE loss, as illustrated in Fig. 4. As shown in the bottom line of Fig. 7, the proposed Anchor Iterative Update loss alleviates the over-saturation problem and generates realistic textures, leading to better generation performance.
4 Experiments
4.1 Implementation Details
Datasets. We train our model on the public available Objaverse [4] dataset which contains around 800k 3D objects. After adopting simple filter leveraging CLIP [27] to remove the objects whose rendered images are not relevant to its name, we have about 350K objects at the end. For each object, we randomly pick two elevation angles and render 32 images with the azimuth evenly distributed in $[0^{\circ},360^{\circ}]$. We then follow the setting of MVDream [31] to render multi-view RGB images for training.
@@ 4.1 Implementation Details
Training. We train our model on 32 NVIDIA V100 GPUs for 30k steps with batchsize 512. The total training process cost about 6 days. We use $256\times 256$ as the image resolution for training. For each batch we randomly sample 4 views as conditional multi-view images and another one view as the target view. The model is initialized from MVDream (the version of stable diffusion 2.1) and the optimizer settings and $\epsilon$-prediction strategy are retained from the previous finetuning except we reduce the learning rate to 1e-5 and use 10 times learning rate for camera encoders’ parameters for faster convergence. As there is no official code for instant3D [13], we reproduce it follows the published paper and use it as our base model to generate the pseudo multi-view images and the coarse initial shape. (More implementation details could be found in the Supplementary Materials).
Figure 5: Qualitative Comparsion bewteen our method with others.
@@
Figure 6: Qualitative comparison between our method with other refinement methods. From left to right, we show input, followed by the initial results from instant3D [13], and culminates with the results derived from StableDiffusion Boost, Zero123-XL [16] Boost, and our multi-view conditioned model.
4.2 Qualitative Comparisons.
@@
We make comparisons with State-Of-the-Art methods on Image-to-3D generation, including Zero123-XL [16] and ImageDream [34]. As shown in Fig. 5, our method generates visually better results with sharper textures, better geometry and highly consistent 3D alignment. Zero-1-to-3 [16] empowers 2D diffusion models with viewpoint conditioning, enabling zero-shot novel view synthesis conditioned on a single-view image. However, limited by the incomplete information provided by a single-view input, it suffers from high uncertainty on the occluded regions, struggling to generate the unseen regions with implausible shapes and chaotic texture. Compared to Zero123, our model benefits from more comprehensive information obtained from the multi-view inputs, migrating the problems of flat geometry and blurry textures and generates high-quality results. ImageDream [34] extends MVDream [31] to start from a given input image, which proposes a new variant of image conditioning. Compared to Zero-1-to-3, ImageDream employs self-attention to establish multi-view correspondence, generating more accurate shapes. However, as shown in Fig. 5, it still generates results with over-saturated colors and over-smooth textures. Comparison to Imagedream, our model benefits from much stronger and more precise constraints provided by the pseudo multi-view inputs, leading to higher generation quality with photo-realistic colors and more geometric details.
@@
We also make comparisons by using other diffusion models at the optimization stage to boost the coarse results, including Stable Diffusion and Zero123-XL [16]. Note we only compare with methods that supports or built on relative camera coordinates. As ImageDream [34] is trained in its own camera canonical space, it could not be used in the boost stage directly. For the fairness, we replace our multi-view conditioned diffusion model with other diffusion models at the boost stage while keeping other settings unchanged. As shown in Fig. 6 and Tab. 1, limited by the ambiguity of text description, Stable Diffusion fails to keep the identity of the initiated coarse object fixed while changing it into another one with different texture and incorrect geometry. Although empowered with single-view image input in the cross-attention to synthesize novel views, Zero123-xl still lacks the ability to generate accurate 3D contents with consistent geometry and detailed texture, especially for the unseen regions. As shown in Fig. 6, it fails to enhance the occluded regions of the initial objects and generates 3D inconsistent texture with even lower qualities. Compared to these methods, our model acquires strong ability of generating highly consistent images from the pesudo multi-view images and provides more precise SDS guidance which can effectively maintain the identity and enhance local details in both geometry and texture of initial generation results.
Table 1: Quantitative Comparisons with image-to-3D methods.
Model QIS CLIP(TX) CLIP(IM) Time
MVDream $27.42\pm 4.92$ $31.42\pm 3.34$ $79.36\pm 4.20$  $\sim 1$h
@@
Basemodel(Instant3D) $23.16\pm 2.53$ $30.61\pm 3.22$ $83.20\pm 5.45$  $\sim 10$s
SD-Boost $22.85\pm 4.37$ $30.26\pm 3.51$ $78.21\pm 10.32$  $\sim 15$min
Zero123-Boost $23.08\pm 3.31$ $29.67\pm 3.09$ $81.73\pm 5.96$  $\sim 15$min
@@
Following ImageDream [34], we adopt three metrics for quantitative comparison of our methods with others, including the Quality-only Inception Score [28] (QIS) and CLIP [27] scores that calculated with text prompts and image prompts, respectively. Among which, QIS evaluates image quality and CLIP scores assess the coherence between the generated models and the prompts. The evaluation dataset consists of 37 high-resolution images, which is generated by SDXL with well-curated prompts. We present the quantitative comparison results in Tab. 1. As shown in the results, our model achieves higher scores on all the three metrics compared to other method, demonstrating the great image conditioned generation ability of our models. As the condition image is synthesized by text prompt, the highest CLIP-Text score also demonstrates the capablity of our model to generates highly text-aligned contents. We also compare the inference time of different methods, which is calculated on a single A100 GPU. Notably, the inference speed of our model is significantly faster than the traditional optimization-based methods like Magic123 [26] and ImageDream [34], while achieving better generation quality.
4.4 Ablation Study.
Number of condition views. In the top row of Fig. 7, we show the effects of our multi-view conditioned diffusion model to synthesize novel views with different number of inputs. As illustrated in the Fig. 7, when there is only one single view input, our model fails to generate view consistently novel-view images. This demonstrates that the diffusion model struggle to acquires accurate 3D information from only one view input, resulting in the inconsistent generation results. However, as the number of the condition views increases, the generation fidelity of our model gradually increases, leading to more consistent results which demonstrates the necessity of our proposed multi-view conditioned diffusion model.
Data augmentation. We compare models trained with/without data augmentation in Tab. 1 and Fig. 7. As shown in the left figure of Fig. 7, model trained without data augment fails to generate 3D consistent results as a result of learning strong condition to certain nearest anchor view while ignoring information from others. In comparison, with the proposed data augmentation strategy, our model learns to fuse the information from all conditional views and generate highly consistent 3D results. Another advantage of our model is to use the condition strength to control the condition strength of each input view, which allows user to strength the influence of high-quality views while weaken the ones with 3D inconsistency or in low-quality. As illustrated in the middle of Fig. 7, by fixing the weights of the first input view while lowing the weights of others, the model learns to generate diverse realistic results without constraining strictly by the input views. By setting the condition strength at different level, our model is capable to deal with inputs in different quality level, which improves the robustness of our model. (Please refer to the supplementary materials for more qualitative results.)
Anchor Iterative Update loss. We adopt Anchor Iterative Update loss to deal with the over-saturation problem of SDS. However, a more straightforward method is to simply use L1 loss from four pesudo multi-view images to constrain the appearance of the object from being awary from the input. However, as illustrated in the bottom line of Fig. 7, L1 loss highly rely on the consistency of the input multi-view images, leading to obvious artifacts with collapsed textures at the inconsistent regions between different views. Instead, Anchor Iterative Update losss, leveraging the iterative render, update and distill strategy, alleviates the over-saturation problem by gradually distilling the clean RGB image refined from the updated rendered anchor images, leading to more robust performance with realistic appearance.
Generation with coarse initialization. In addition, we evaluated the impact of utilizing the coarse reconstruction outcomes obtained from a larger reconstruction model as an initialization step. As illustrates in Fig. 8, when provided with pseudo multi-view inputs, our model is capable of producing high-quality 3D contents from scratch. However, this process requires approximately ×$\times 5$ time (1.2h) for our model to achieve comparable results to those generated using the coarse reconstruction outputs as initialization. This result emphasizes the effectiveness of our overall approach.
5 Limitation.
6 Conclusion
We present Magic-Boost, a multi-view conditioned diffusion model which takes pseudo generated multi-view images as input, implicitly encodes 3D information from different views, and capable of synthesising highly consistent novel view images and providing precise SDS guidance during the optimization process. Extensive experiments demonstrate our model greatly enhances the generation quality of coarse input and generates high-quality 3D assets with detailed geometry and realistic texture within a short time period.
Appendix 0.A Appendix
In this Appendix, we provide additional implementation details in Sec. 0.A.1, evaluation on the novel view synthesis in Sec. 0.A.2, additional quliatative comparisons in Sec. 0.A.3. We present limitations and future work in Sec. 0.A.4.
0.A.1 Additional Implementation Details.
Network Structure. We build our network on the Stable Diffusion backbone, while we make several changes on the structure: 1) We incorporate a frozen CLIP pre-trained Vision Transformer [27] as global feature extractor. 2) We adopt the same denoising U-Net with shared weights to extract local features at fixed time step 00. 3) We extend the original 2D self attention layer into 3D by concatenating keys and values from different views to facilitate information propagation. 4) We encode the camera pose with a two layer MLP, which is then added to time embedding as residuals. The proposed control label is embedded and added to the time embedding in the same way.
0.A.2 Evaluation on Novel View Synthesis.
@@ Appendix 0.A Appendix
We employ PSNR, SSIM [36], LPIPS [42] to evaluate the performance of our model on novel view synthesis task on the Google Scaned Dataset (GSO) [6]. In details, we use the same 30 objects chosen by SyncDreamer [17] and render 16 views with uniformly distributed camera poses and environment lighting for each object. To ensure a fair and efficient evaluation process, the first render view is selected as the input image for each baseline and our methods. For multi-view inputs of our method, we select views orthogonal to the first input view. As shown in Tab. 2, benefit from the more comprehensive information provided by multi-view inputs, our model significantly outperforms previous methods that rely on only single-view image as input and synthesize novel views with much higher fidelity. We provide more novel view synthesis results in Fig. 9.
@@
Table 2: Quantitative Comparisons in Novel view synthesis. We report PSNR, SSIM [36], LPIPS [42] on the GSO [6] dataset.
@@
0.A.3 More Qualitative Results.
We provide more qualitative results in this Section. Fig. 10 and Fig. 11 shows more results of our Image-to-3D generation results. In this Figure, we show from left to right: the input image, the generated multi-view images, the coarse output from Instant3D and the boosted results of our method. In Fig. 12 and Fig. 13, we show results of text-to-3D generation. In text-to-3D setting, we directly generates the multi-view images from the given text input. In Fig. 12 and Fig. 13. from left to right, we show the input text, the generated multi-view images, the coarse output and the boosted results of our method. In Fig. 14, we show more qualitative comparisons of our method with other possible boost method, including using Stable Diffusion to boost the results (Sd-boost) and using Zero123-XL [16] to boost the results (image-boost). As shown in Fig. 14, our model provides more precise SDS guidance which can effectively maintain the identity and enhance local details in both geometry and texture of initial generation results. We also illustrate the effects of the control label in Fig. 15, while using high control strength the model highly rely on the consistency of the input multi-view images, leading to chaotic texture in the inconsistent regions. When we reduce the control strength of the back view here, the model generates smoother results with better visual quality.
@@
Figure 10: More results of Image conditioned generation. From left to right: the input image, the generated multi-view images, the coarse output from Instant3D and the boosted results of our method.
Figure 11: More results of Image conditioned generation. From left to right: the input image, the generated multi-view images, the coarse output from Instant3D and the boosted results of our method.
Figure 12: More results of text conditioned generation. From left to right, we show the input text, the generated multi-view images, the coarse output and the boosted results of our method.
Figure 13: More results of text conditioned generation. From left to right, we show the input text, the generated multi-view images, the coarse output and the boosted results of our method.
Figure 14: More comparision results of our boost method with ours, including sd-boost (Stable Diffusion) and image-boost (Zero123-XL [16]). As shown in the figure, our model can effectively maintain the identity and enhance local details in both geometry and texture of coarse generation results.
Figure 15: Illustration of the effects of control label. While using high control strength the model highly rely on the consistency of the input multi-view images, leading to chaotic texture in the inconsistent regions. When we reduce the control strength of the back view here, the model generates smoother results with better visual quality.
Figure 16: Illustration of a failure case. the model fails to generated correct results on the tail of the horse due to the seriously 3D inconsistent on the generated multi-view images. Exploring 2D multi-view diffusion models with better generation quality would further improve our model’s performance.
0.A.4 Failure Cases.
The use of pseudo 3D multi-view images synthesized by the finetined 2D multi-view diffusion model as inputs imposes an unavoidable influence on our model’s performance. Despite implementing data augmentation during training to emulate 3D inconsistencies and attempting to bridge the domain gap, this approach still results in inaccuracies for faulty structures. For instance, as shown in Fig. 16, the model fails to generated correct results on the tail of the horse due to the seriously 3D inconsistent on the generated multi-view images. Exploring 2D multi-view diffusion models with better generation quality would further improve our model’s performance.

### v3

@@ III-C Refinement with SDS Optimization
As shown in Figure 2, We build our test pipeline with Instant3D [1], a two-stage feed-forward generation method which firstly generates four multi-view images with finetuned 2D diffusion networks and then lift the multi-view images to 3D model utilizing large reconstruction model. However, due to limited resolution and local inconsistency, the generated results are still in low quality without detailed texture and complicated geometry. Leveraging the generated pseudo multi-view as input, we propose to adopt SDS optimization with small noise level to further enhance the coarse generation results. Benefit from the high-fidelity geometry and texture information provided by the pesudo multi-view inputs, our model is capable of generating highly consistent novel views and providing precise SDS guidance to enhance the coarse generated results within a short time period ($\sim 15$min).
@@
We further introduce an Anchor Iterative Update loss to alleviate the over-saturation problem of SDS optimization. Specifically, we draw inspiration from the recent image editing methods [41, 42] that edit the NeRF by alternatively rendering dataset image, updating the dataset and then supervising the NeRF reconstruction with the updated dataset images. In our generation task, we could also regard the input pesudo multi-view inputs as our initial anchor datasets and adopt a similar update strategy by first rendering anchor view image, perturbing the image with random noise and then apply a multi-step denoising process with the proposed multi-view condition diffusion model to refine the anchor images. As the denoising process is guided by the multi-view inputs themselves, simply adopt such as process leading to minor refinement on the input anchor images. To address this problem, while updating certain anchor view $v_{1}$, we drop it in the conditional inputs, and denoise leveraging other views $v_{0},v_{2},v_{3}$. We found this works well and achieve accurate refinement on the local details of the anchor image, while preserving the 3D consistency. The refined anchor image is then used to supervise the generation with MSE loss, as illustrated in Figure 4. As shown in the bottom line of Figure 7, the proposed Anchor Iterative Update loss alleviates the over-saturation problem and generates realistic textures, leading to better generation performance.
IV Experiments.
IV-A Implementation Details
@@
Training. We train our model on 32 NVIDIA V100 GPUs for 30k steps with batchsize 512. The total training process cost about 6 days. We use $256\times 256$ as the image resolution for training. For each batch we randomly sample 4 views as conditional multi-view images and another one view as the target view. The model is initialized from MVDream (the version of stable diffusion 2.1) and the optimizer settings and $\epsilon$-prediction strategy are retained from the previous setting for finetuning except we reduce the learning rate to 1e-5 and use 10 times learning rate for camera encoders’ parameters for faster convergence. We train our model on the public available Objaverse [28] dataset following the setting of MVDream [40]. Please refer to Appendix for more implementary details.
Testing. We adopt Instant3D [1], one of the SOTA 3D generation methods that involves multi-view generation and fast reconstruction as our baseline model to provide the multi-view priors and coarse reconstruction results. As there is no official code for instant3D [1], we reproduce it follows the published paper. However, as discussed in Section IV-D, our model could also be plugged into other models capable of providing pseudo multi-view priors, such as InstantMesh [14], LGM [15] and etc.
Figure 5: Qualitative Comparison between our method with Imagedream [38] and base sparse-view reconstruction model [1]. SVR denotes Sparse-View Reconstruction.
@@
Figure 6: Qualitative comparison between our method with other refinement methods. From left to right, we show input, followed by the reconstruction results from sparse view reconstruction models [1], and the results refined by StableDiffusion [11], Zero123-XL [13], and Ours.
IV-B Evaluation on Image-to-3D Generation
IV-B1 Qualitative Comparisons.
@@ IV-B Evaluation on Image-to-3D Generation
We make comparisons with State-Of-the-ArT methods on Image-to-3D generation task, ImageDream [38]. As shown in Figure 5, our method generates visually better results with sharper textures, better geometry and highly consistent 3D alignment. ImageDream [38] extends MVDream [40] to start from a given input image, which proposes a new variant of image conditioning. However, limited by the incomplete information provided by a single-view input, it suffers from high uncertainty on the occluded regions, struggling to generate the unseen regions with implausible shapes and chaotic texture. In contrast, our model benefits from much stronger and more precise 3d prior provided by the pseudo multi-view inputs, leading to higher generation quality with photo-realistic colors and more geometric details.
@@
We also make comparisons by using other diffusion models at the optimization stage to refine the sparse view reconstruction results, including Stable Diffusion [11] and Zero123-xl [13]. As ImageDream [38] is trained in its own camera canonical space, it could not be used in the refinement stage directly. For the fairness, we replace our multi-view conditioned diffusion model with other diffusion models at the refinement stage while keeping other settings unchanged. As shown in Figure 6 and Table I, limited by the ambiguity of text description, Stable Diffusion fails to keep the identity of the initiated coarse object fixed while changing it into another one with different texture and incorrect geometry. Although empowered with single-view image input in the cross-attention to synthesize novel views, Zero123-xl still lacks the ability to generate accurate 3D contents with consistent geometry and detailed texture, especially for the unseen regions. Compared to these methods, our model acquires strong ability of generating highly consistent images from the pseudo multi-view images and provides more precise SDS guidance which can effectively maintain the identity and enhance local details in both geometry and texture of initial generation results.
TABLE I: Quantitative Comparisons with image-to-3D methods.
[…]
@@
Baseline model [1] $23.16\pm 2.53$ $30.61\pm 3.22$ $83.20\pm 5.45$ $\sim 10$s
SD-Refine $23.85\pm 4.37$ $32.26\pm 3.51$ $78.21\pm 10.32$ $\sim 15$min
Zero123-Refine [13] $24.08\pm 3.31$ $31.67\pm 3.09$ $84.73\pm 5.96$ $\sim 15$min
@@
Following ImageDream [38], we adopt three metrics for quantitative comparison of our methods with others, including the Quality-only Inception Score [43] (QIS) and CLIP [39] scores that calculated with text prompts and image prompts, respectively. Among which, QIS evaluates image quality and CLIP scores assess the coherence between the generated models and the prompts. The evaluation dataset consists of 37 high-resolution images, which is generated by SDXL with well-curated prompts. We present the quantitative comparison results in Table I. As shown in the results, our model achieves higher scores on all the three metrics compared to other method, demonstrating the great image conditioned generation ability of our models. As the condition image is synthesized by text prompt, the highest CLIP-Text score also demonstrates the capablity of our model to generates highly text-aligned contents. We also compare the inference time of different methods, which is calculated on a single A100 GPU. Notably, the inference speed of our model is significantly faster than the traditional optimization-based methods like Magic123 [8] and ImageDream [38], while achieving better generation quality.
IV-C Evaluation on Novel View Synthesis
@@
We employ PSNR, SSIM [44], LPIPS [45] to evaluate the performance of our model on novel view synthesis task on the Google Scaned Dataset (GSO) [46]. In details, we use the same 30 objects chosen by SyncDreamer [2] and render 16 views with uniformly distributed camera poses and environment lighting for each object. To ensure a fair and efficient evaluation process, the first render view is selected as the input image for each baseline and our methods. For multi-view inputs of our method, we select views orthogonal to the first input view. As shown in II, benefit from the more comprehensive information provided by multi-view inputs, our model significantly outperforms previous methods that rely on only single-view image as input and synthesize novel views with much higher fidelity. We provide more novel view synthesis results in Figure 10.
@@
TABLE II: Quantitative Comparisons in Novel view synthesis. We report PSNR, SSIM [44], LPIPS [45] on the GSO [46] dataset.
@@
IV-D Ablation Study.
Number of condition views. We show the effects of our multi-view conditioned diffusion model to synthesize novel views with different number of conditional views in Figure 7 and Table II. When there is only one single view input, our model fails to generate consistent novel-view images. This demonstrates that the diffusion model struggle to acquires accurate 3D information from only one view input, resulting in the inconsistent generation results. However, as the number of the condition views increases, the generation fidelity of our model gradually increases, leading to more consistent results which demonstrates the necessity of our proposed multi-view conditioned diffusion model.
Data augmentation. We compare models trained with/without data augmentation in Table I and Figure 7. As shown in Figure 7, the Model trained without data augment fails to generate 3D consistent results due to the domain gap between training stage and inference stage. In comparison, with the proposed data augmentation strategy, our model learns to better correct the inconsistency in the conditional views and generate highly consistent 3D results. Figure 7 also illustrates the effect of the proposed control label. By fixing the weights of the first input view while lowing the weights of others, the model learns to generate diverse realistic results without constraining strictly by the input views. By setting the condition strength at different level, our model is capable to deal with inputs in different quality level, which improves the robustness of our model.
Anchor Iterative Update loss. We adopt Anchor Iterative Update loss to deal with the over-saturation problem of SDS. A more straightforward method is to simply use L1 loss from four pesudo multi-view images to constrain the appearance of the object from being awary from the input. However, as illustrated in the bottom line of 7, L1 loss highly rely on the consistency of the input multi-view images, leading to obvious artifacts with collapsed textures at the inconsistent regions between different views. Instead, Anchor Iterative Update losss, leveraging the iterative render, update and distill strategy, alleviates the over-saturation problem by gradually distilling the clean RGB image refined from the updated rendered anchor images, leading to more robust performance with realistic appearance.
Figure 8: Our model could also be plugged into other models capable of providing pseudo multi-view priors. In this figure we show boost results of our model pluged into InstantMesh [14]. (See appendix for more visual results).
Plug in different 3D generation models. Although we adopt Instant3D [1] as our baseline model in our experiments. Our model could also be plugged into other models capable of providing pseudo multi-view priors, such as InstantMesh [14], LGM [15] and etc. Figure 8 shows an example of adopting InstantMesh [14] as our base model and demonstrates the generalization ability of our model for different 3D generation methods. (See appendix for more visual results).
Generation with coarse initialization. In addition, we evaluated the impact of utilizing the coarse reconstruction outcomes obtained from a larger reconstruction model as an initialization step. As illustrates in Figure 9, when provided with pseudo multi-view inputs, our model is capable of producing high-quality 3D contents from scratch. However, this process requires approximately ×$\times 5$ time (1.2h) for our model to achieve comparable results to those generated using the coarse reconstruction outputs as initialization. This result emphasizes the effectiveness of our overall approach.
V Limitation.
VI Conclusion
We present Magic-Boost, a multi-view conditioned diffusion model which takes pseudo generated multi-view images as input, capable of synthesising highly consistent novel view images and providing precise SDS guidance during the optimization process. Extensive experiments demonstrate our model greatly enhances the generation quality of coarse input and generates high-quality 3D assets with detailed geometry and realistic texture within a short time period.
[Additional Implementation Details] Network Structure. We build our network on the Stable Diffusion backbone [11], while we make several changes on the structure: 1) We incorporate a frozen CLIP pre-trained Vision Transformer [39] as global feature extractor. 2) We adopt the same denoising U-Net with shared weights to extract local features at fixed time step 00. 3) We extend the original 2D self attention layer into 3D by concatenating keys and values from different views to facilitate information propagation. 4) We encode the camera pose with a two layer MLP, which is then added to time embedding as residuals. The proposed control label is embedded and added to the time embedding in the same way.
Appendix A More Qualitative Results.
Please refer to our project page: project page for video results and more qualitative comparisons and our GitHub for code: code. Figure 11 shows more results our method based on InstantmMesh [14].
@@ VI Conclusion
Figure 11: More results of Image conditioned generation based on InstantMesh [14] From left to right: the input image, the pseudo multi-view images, the sparse view reconstruction results and the boosted results of our method.